# GoalRank: Group-Relative Optimization for a Large Ranking Model

### v1

@@ 5 Conclusion
For the theoretical results, detailed derivations are provided in Appendix B. For the empirical studies, we will release the implementation and training code at https://github.com/Kaike-Zhang/GoalRank to ensure reproducibility.
Appendix A Related Work
The ranking task in recommender systems can be viewed as an $N{\!\rightarrow\!}L$ list-generation problem: given $N$ candidate items, the model outputs an ordered list of length $L$. Since the search space $\mathrm{P}(N,L)=\tfrac{N!}{(N-L)!}$ is combinatorially large, a series of paradigms have been developed to approximate this process.
One-stage generator-only methods. Early approaches adopt a single generator that directly produces ranking lists by scoring items and arranging them greedily (Zhuang et al., 2018; Ai et al., 2018; Pei et al., 2019a; Gong et al., 2022; Liu et al., 2023; Feng et al., 2021b; Xi et al., 2022; Pei et al., 2019b). Representative models include the GRU-based DLCM (Ai et al., 2018) and Transformer-based PRM (Pei et al., 2019a), which refine scores using listwise features. While simple and efficient, these methods largely ignore interdependencies among items or suffer from inconsistencies with the initial candidate conditioning, leading to suboptimal results.
Two-stage Generator–Evaluator (G–E) methods. To address these issues, subsequent works propose a two-stage Generator–Evaluator paradigm (Chen et al., 2022; Shi et al., 2023; Xi et al., 2024; Lin et al., 2024; Ren et al., 2024b; Wang et al., 2025b). In this framework, the generator proposes multiple candidate lists, and the evaluator selects the best one by estimating listwise utility. Compared with heuristic permutation strategies, generative approaches are more effective at exploring the exponentially large space while incorporating user preferences, and thus have become the mainstream solution.
Multi-Generator–Evaluator extensions. More recent studies extend the G–E paradigm by employing multiple generators to enlarge the candidate-list space and enhance ranking performance (Yang et al., 2025). However, the improvement is substantial only when the number of generators is small, and quickly diminishes as more generators are added. This indicates that merely scaling the number of generators is insufficient, motivating the search for fundamentally stronger rankers.
Other directions. Beyond these paradigms, recent work has explored the use of large language models (Ren et al., 2024a; Gao et al., 2024; Wu et al., 2024; Gao et al., 2025; Ren et al., 2025; Liu et al., 2025) and reinforcement learning (Feng et al., 2021c; Wang et al., 2024; 2025c; Wei et al., 2020) to improve ranking. These methods incorporate textual side information or treat ranking as a critic-based paradigm, decomposing listwise value estimation to better align recommendations with user preferences.
Appendix B Proof of Theorem 1
The detailed proof will be made publicly available after the paper is accepted for publication.
Appendix C Reward Model Training

### v2

@@ 5 Conclusion
Reproducibility Statement. For the theoretical results, detailed derivations are provided in Appendix A. For the empirical studies, we will release the implementation and training code at https://anonymous.4open.science/r/GoalRank to ensure reproducibility.
Ethics Statement. This work aims to improve the ranking stage of recommender systems to enhance user satisfaction. It does not raise any specific ethical concerns.
Appendix A Proof of Theorem 1
Notation.
Recall the space of the ranking list in Section 2. For simplicity and clarity, in this section, we omit the $u\in\mathcal{U}$ in the expressions of the following symbols. The space of ranking list $\mathcal{L}$ is:
$\mathcal{L}:=\bigl\{\,l=(v_{1},\dots,v_{L})\in\mathcal{V}^{L}\ \big|\ v_{i}\neq v_{j}\ (i\neq j)\,\bigr\},\quad|\mathcal{L}|=\mathrm{P}(N,L):=\frac{N!}{(N-L)!}.$
Then the state set during the generation process can be expressed as:
$\mathcal{S}:=\bigl\{(i,l_{<i})\ \big|\ i\in[L],\ l_{<i}\in\mathcal{V}^{i-1},\ v_{a}\neq v_{b}\ (a\neq b)\bigr\}.$
and corresponding actions at state $(i,l_{<i})$:
$\mathcal{A}(i,l_{<i}):=\mathcal{V}\setminus\{v_{1},\dots,v_{i-1}\},\quad|\mathcal{A}(i,l_{<i})|=N-(i-1).$
Given a score vector $a\in\mathbb{R}^{\mathcal{V}}$ and a feasible set $\mathcal{A}\subseteq\mathcal{V}$, the masked softmax is
$\bigl(\mathrm{softmax}_{\mathcal{A}}(a)\bigr)_{j}\;:=\;\begin{cases}\dfrac{e^{a_{j}}}{\sum_{t\in\mathcal{A}}e^{a_{t}}},&j\in\mathcal{A},\\[8.0pt] 0,&j\notin\mathcal{A}.\end{cases}$
For a finite index set $\mathcal{I}$ and $x\in\mathbb{R}^{\mathcal{I}}$, write $\|x\|_{\infty;\mathcal{I}}:=\max_{i\in\mathcal{I}}|x_{i}|$. We use $\mathrm{KL}(\cdot\|\cdot)$ for Kullback–Leibler divergence.
Lemma 1.
Let $k\in\mathbb{N}$ and $g_{m,i}\in\mathcal{G}_{m}$ ($i=1,\dots,k$) be $k$ generators. Assume each $g_{m,i}$ is locally Lipschitz in its parameter $\theta_{i}\in[A,B]^{d_{m}}$, and the masking set $\mathcal{A}(t,l_{<t})$ depends only on $(t,l_{<t})$ (not on parameters). For mixing weights $\bm{\omega}=[\omega_{i}]_{i=1}^{k}\in\Delta^{k-1}$, define
$\Phi:\ \Theta_{k}\to\Delta^{|\mathcal{L}|-1},\qquad\Theta_{k}:=\big([A,B]^{d_{m}}\big)^{k}\times\Delta^{k-1},$ (7)
where
$\Phi(\bm{\theta}_{1:k},\bm{\omega})=\sum_{i=1}^{k}\omega_{i}\,\pi_{g_{m,i}},$
where $\pi_{g_{m,i}}=\big(\pi_{g_{m,i}}(l)\big)_{l\in\mathcal{L}}$ is the masked-softmax autoregressive policy induced by $g_{m,i}$. Then $\mathcal{C}_{m}^{k}:=\mathrm{im}(\Phi)$ satisfies:
  1. 1.
    $\mathcal{C}_{m}^{k}$ is compact;
  2. 2.
    $\dim_{\mathrm{Haus}}\big(\mathcal{C}_{m}^{k}\big)\leq\min\big\{kd_{m}+(k-1),\ |\mathcal{L}|-1\big\}$.
Proof.
Compactness. Each parameter domain $A,B]^{d_{m}}$ is compact, and finite Cartesian products preserve compactness; the simplex $\Delta^{k-1}$ is also compact. Hence $\Theta_{k}$ is compact. For a fixed list $l\in\mathcal{L}$ and position $t$, given the smoothness of softmax and $g_{m,i}$ is $C^{1}$, the masked-softmax distribution
$\pi_{g_{m,i}}(l)=\prod_{t=1}^{L}\frac{\exp\big(z^{(i)}_{t,l_{t}}(l_{<t})\big)}{\sum_{j\in\mathcal{A}(t,l_{<t})}\exp\big(z^{(i)}_{t,j}(l_{<t})\big)},\qquad z^{(i)}_{t,\cdot}=g_{m,i}(t,l_{<t},\cdot;\theta_{i}),$
is $C^{1}$ in $\theta_{i}$, as masking only discards coordinates independent of parameters.
Since $\Phi$ is a convex combination of such policies, it is $C^{1}$ in $(\bm{\theta}_{1:k},\bm{\omega})$, hence continuous. By continuity of $\Phi$ on the compact domain $\Theta_{k}$, its image $\mathcal{C}_{m}^{k}$ is compact.
Dimension bound. On $\Theta_{k}$ (compact), each map $(\theta_{i},\bm{\omega})\mapsto\omega_{i}\pi_{g_{m,i}}$ is Lipschitz: $\theta_{i}\mapsto z^{(i)}$ is Lipschitz; composition with masked-softmax is Lipschitz (bounded Jacobian on the relevant compact image); the product over $t$ and convex mixing in $\bm{\omega}$ preserve Lipschitzness since all factors are uniformly bounded. Thus $\Phi$ is Lipschitz on $\Theta_{k}$. Lipschitz maps do not increase Hausdorff dimension, yielding
$\dim_{\mathrm{Haus}}(\mathrm{im}(\Phi))\leq\dim_{\mathrm{Haus}}(\Theta_{k})=kd_{m}+(k-1),$
and trivially $\dim_{\mathrm{Haus}}(\mathrm{im}(\Phi))\leq|\mathcal{L}|-1$, giving the stated minimum. ∎
Theorem 2.
Let $\mathcal{V}$ be the candidate set with $N:=|\mathcal{V}|$, and let $\mathcal{L}$ be the set of length-$L$ lists without repetition, so $d:=|\mathcal{L}|-1=\mathrm{P}(N,L)-1$. Let $\mathcal{C}_{m}^{k}(\alpha,\beta)\subset\Delta^{d}$ be the $k$-mixture $(\alpha,\beta)$-bounded policy space. Assume Lemma 1 holds with
$\dim_{\mathrm{Haus}}\!\bigl(\mathcal{C}_{m}^{k}(\alpha,\beta)\bigr)\ \leq\ r:=k\,d_{m}+(k-1)\qquad\text{and}\qquad r<d.$
Then for Lebesgue-almost every fully supported target policy $\pi^{*}\in\mathrm{int}(\Delta^{d})$,
$\inf_{\pi\in\mathcal{C}_{m}^{k}(\alpha,\beta)}\mathrm{KL}\!\bigl(\pi^{*}\|\pi\bigr)\;>\;0.$ (8)
Proof.
By Lemma 1, $\mathcal{C}_{m}^{k}(\alpha,\beta)$ is compact and $\dim_{\mathrm{Haus}}(\mathcal{C}_{m}^{k}(\alpha,\beta))\leq r<d=\dim(\mathrm{aff}(\Delta^{d}))$. Hence the $d$-dimensional relative Lebesgue measure (equivalently, the $d$-dimensional Hausdorff measure) of $\mathcal{C}_{m}^{k}(\alpha,\beta)$ inside $\mathrm{aff}(\Delta^{d})$ is zero. Therefore, for Lebesgue-a.e. $\pi^{*}\in\mathrm{int}(\Delta^{d})$ we have $\pi^{*}\notin\mathcal{C}_{m}^{k}(\alpha,\beta)$.
Fix such a $\pi^{*}$ with $\pi^{*}\succ 0$. Every $\pi\in\mathcal{C}_{m}^{k}(\alpha,\beta)$ assigns strictly positive probability to each $l\in\mathcal{L}$, so the map
$\Psi(\pi)\ :=\ \mathrm{KL}\!\bigl(\pi^{*}\|\pi\bigr)\ =\ \sum_{l\in\mathcal{L}}\pi^{*}(l)\,\log\frac{\pi^{*}(l)}{\pi(l)}$
is finite and continuous on the compact set $\mathcal{C}_{m}^{k}(\alpha,\beta)$; thus the minimum
$\delta\ :=\ \min_{\pi\in\mathcal{C}_{m}^{k}(\alpha,\beta)}\mathrm{KL}\!\bigl(\pi^{*}\|\pi\bigr)$
is attained. Since $\mathrm{KL}(\pi^{*}\|\pi)=0$ iff $\pi=\pi^{*}$ and $\pi^{*}\notin\mathcal{C}_{m}^{k}(\alpha,\beta)$, we must have $\delta>0$, proving Equation 8. ∎
Classical universal approximation results (Cybenko, 1989; Hornik et al., 1989; Hornik, 1991; Leshno et al., 1993; Sonoda and Murata, 2017) yield the following lemma. For clarity and generality, we adopt MLP-based generators as the foundational model class. The goal is to show that even the most basic architecture—the MLP—already has sufficient expressive power to approximate the target policy. Universal approximation for Transformers (and related architectures) in sequence modeling is also known; see Yun et al. (2019); Augustine (2024).
Lemma 2.
Let $K\subset\mathbb{R}^{n}$ be compact and let $F:K\to\mathbb{R}^{m}$ be continuous. Let $\phi:\mathbb{R}\to\mathbb{R}$ be any continuous activation function that is not a polynomial on any interval. [^3] This covers common activations such as sigmoid, $\tanh$, ReLU, leaky-ReLU, ELU, and softplus. Then for every $\eta>0$ there exists a fixed-depth (e.g., one hidden layer), arbitrarily wide MLP $h_{\Theta}:K\to\mathbb{R}^{m}$ with activation $\phi$ such that
$\|h_{\Theta}-F\|_{\infty;K}\;<\;\eta.$
Consequently, for any finite Borel measure $\mu$ on $K$ and any $1\leq p<\infty$, we also have $\|h_{\Theta}-F\|_{L^{p}(K,\mu)}<\eta$.
Remark 1 (Finite domains).
On a finite domain $K=\{x_{1},\dots,x_{T}\}$, continuity is automatic and
$\|h_{\Theta}-F\|_{\infty;K}=\max_{t\in[T]}\|h_{\Theta}(x_{t})-F(x_{t})\|_{\infty},$
which coincides with the maximum pointwise error.
To adapt Lemma 2 to the ranking setting, we encode the autoregressive state–action tuples into a compact Euclidean set.
Corollary 1.
Let
$\rho:\mathcal{D}\to\mathbb{R},\qquad\mathcal{D}:=\bigl\{(i,l_{<i},j)\ :\ 1\leq i\leq L,\ j\in\mathcal{A}(i,l_{<i})\bigr\},$
be defined by
$\rho(i,l_{<i},j)\ :=\ \log\pi^{*}(l_{i}=j\mid l_{<i}),$
where $\pi^{*}$ is a fully supported target autoregressive policy on the feasible sets $\mathcal{A}(i,l_{<i})$. Fix any injective encoding $\psi:\mathcal{D}\hookrightarrow K\subset\mathbb{R}^{d}$ with $K$ compact. Then for any $\sigma>0$ there exists an MLP $h_{\Theta}:K\to\mathbb{R}$ such that
$\|\,h_{\Theta}\circ\psi-\rho\,\|_{\infty;\mathcal{D}}\ \leq\ \sigma.$
Proof.
The index set $\mathcal{D}$ is finite (since $N,L<\infty$), hence any function defined on $\psi(\mathcal{D})\subset K$ is continuous w.r.t. the subspace topology. Apply Lemma 2 with $m=1$ to obtain $h_{\Theta}$ with the desired uniform bound. ∎
Lemma 3.
Let $p(j)\propto e^{\rho_{j}}$ on a finite set and $q(j)\propto e^{\rho_{j}+\varepsilon_{j}}$ with $|\varepsilon_{j}|\leq\sigma$ for all $j$. Then for all $j$,
$e^{-2\sigma}\ \leq\ \frac{q(j)}{p(j)}\ \leq\ e^{2\sigma},$
and in particular
$\mathrm{KL}(p\|q)\ \leq\ 2\sigma,\qquad\|p-q\|_{1}\ \leq\ e^{2\sigma}-1.$
Moreover, for an autoregressive policy over length-$L$ lists,
$\mathrm{KL}\bigl(\pi^{*}\|\tilde{\pi}\bigr)\ =\ \sum_{i=1}^{L}\mathbb{E}_{l_{<i}\sim\pi^{*}}\!\Bigl[\,\mathrm{KL}\bigl(p_{i}(\cdot\mid l_{<i})\|q_{i}(\cdot\mid l_{<i})\bigr)\Bigr]\ \leq\ 2L\sigma.$
Fact 1.
Fix a feasible set $\mathcal{A}$. Let $p=\mathrm{softmax}_{\mathcal{A}}(a)$ and $q=\mathrm{softmax}_{\mathcal{A}}(b)$ with $a,b\in\mathbb{R}^{\mathcal{V}}$. If $\|a-b\|_{\infty;\mathcal{A}}\leq\sigma$, then
$\mathrm{KL}(p\|q)\ \leq\ 2\sigma.$
Proof.
Write $\varepsilon_{j}:=b_{j}-a_{j}$ for $j\in\mathcal{A}$, so $|\varepsilon_{j}|\leq\sigma$. Then
$\frac{q_{j}}{p_{j}}=\frac{e^{b_{j}}/\sum_{t\in\mathcal{A}}e^{b_{t}}}{e^{a_{j}}/\sum_{t\in\mathcal{A}}e^{a_{t}}}=\frac{e^{\varepsilon_{j}}}{\sum_{t\in\mathcal{A}}p_{t}e^{\varepsilon_{t}}}\in\Bigl[\frac{e^{-\sigma}}{e^{\sigma}},\,\frac{e^{\sigma}}{e^{-\sigma}}\Bigr]=\bigl[e^{-2\sigma},\,e^{2\sigma}\bigr].$
Hence $\log\frac{p_{j}}{q_{j}}\leq 2\sigma$ for all $j\in\mathcal{A}$, and
$\mathrm{KL}(p\|q)=\sum_{j\in\mathcal{A}}p_{j}\log\frac{p_{j}}{q_{j}}\ \leq\ \sum_{j\in\mathcal{A}}p_{j}\cdot 2\sigma\ =\ 2\sigma.$
Fact 2.
Let $\pi^{*}$ and $\pi$ be autoregressive list policies on $\mathcal{V}$ of length $L$ with conditionals supported on $\mathcal{A}(i,l_{<i})$. Then
$\mathrm{KL}(\pi^{*}\|\pi)=\sum_{i=1}^{L}\mathbb{E}_{l_{<i}\sim\pi^{*}}\!\Big[\mathrm{KL}\big(\pi^{*}(\cdot\mid l_{<i})\,\|\,\pi(\cdot\mid l_{<i})\big)\Big].$
Proof.
By the chain rule,
$\log\frac{\pi^{*}(l_{1},\dots,l_{L})}{\pi(l_{1},\dots,l_{L})}=\sum_{i=1}^{L}\log\frac{\pi^{*}(l_{i}\mid l_{<i})}{\pi(l_{i}\mid l_{<i})}.$
Taking expectation w.r.t. $\pi^{*}$ yields
$\mathrm{KL}(\pi^{*}\|\pi)=\sum_{i=1}^{L}\mathbb{E}_{l_{<i}\sim\pi^{*}}\,\mathbb{E}_{l_{i}\sim\pi^{*}(\cdot\mid l_{<i})}\!\Bigl[\log\frac{\pi^{*}(l_{i}\mid l_{<i})}{\pi(l_{i}\mid l_{<i})}\Bigr],$
which is the stated form because the inner expectation equals $\mathrm{KL}\big(\pi^{*}(\cdot\mid l_{<i})\|\pi(\cdot\mid l_{<i})\big)$. ∎
Theorem 3.
Let $\mathcal{G}_{\mathrm{MLP}}$ be the class of (fixed-depth, arbitrary-width) MLP generators that, given $(i,l_{<i})$, produce logits $g(i,l_{<i},\cdot)\in\mathbb{R}^{\mathcal{V}}$. Define the induced policy class
$\mathcal{F}_{\mathrm{MLP}}:=\bigl\{\ \pi_{\Theta}\ :\ \pi_{\Theta}(\,\cdot\mid l_{<i})=\mathrm{softmax}_{\mathcal{A}(i,l_{<i})}\bigl(g_{\Theta}(i,l_{<i},\cdot)\bigr)\ \bigr\}.$
Then for every $\varepsilon>0$ there exists a width $W(\varepsilon,N,L)$ and parameters $\Theta$ such that
$\mathrm{KL}(\pi^{*}\|\pi_{\Theta})\ \leq\ \varepsilon,\qquad\text{hence}\qquad\lim_{W\to\infty}\ \mathcal{E}(\mathcal{F}_{\mathrm{MLP}})\ =\ 0,$
where $\mathcal{E}(\mathcal{F}):=\inf_{\pi\in\mathcal{F}}\mathrm{KL}(\pi^{*}\|\pi)$.
Proof.
Define target logits on the effective domain $\mathcal{D}$ by $\rho(i,l_{<i},j):=\log\pi^{*}(l_{i}=j\mid l_{<i})$ for $j\in\mathcal{A}(i,l_{<i})$. Fix an injective encoding $\psi:\mathcal{D}\hookrightarrow K\subset\mathbb{R}^{d}$ into a compact $K$.
By Corollary 1, for any $\sigma>0$ there exists an MLP $h_{\Theta}:K\to\mathbb{R}$ such that $\|h_{\Theta}\circ\psi-\rho\|_{\infty;\mathcal{D}}\leq\sigma$. Use $g_{\Theta}(i,l_{<i},j):=h_{\Theta}(\psi(i,l_{<i},j))$ for $j\in\mathcal{V}$. Define the policy
$\pi_{\Theta}(l_{i}=j\mid l_{<i}):=\mathrm{softmax}_{\mathcal{A}(i,l_{<i})}\bigl(g_{\Theta}(i,l_{<i},\cdot)\bigr)_{j}.$
For each prefix $l_{<i}$, Fact 1 with $a=\rho(i,l_{<i},\cdot)$, $b=g_{\Theta}(i,l_{<i},\cdot)$ yields
$\mathrm{KL}\!\left(\pi^{*}(\cdot\mid l_{<i})\ \|\ \pi_{\Theta}(\cdot\mid l_{<i})\right)\ \leq\ 2\sigma.$
Applying Fact 2 gives
$\mathrm{KL}(\pi^{*}\|\pi_{\Theta})\ =\ \sum_{i=1}^{L}\mathbb{E}_{l_{<i}\sim\pi^{*}}\!\left[\mathrm{KL}\bigl(\pi^{*}(\cdot\mid l_{<i})\|\pi_{\Theta}(\cdot\mid l_{<i})\bigr)\right]\ \leq\ 2L\sigma.$
Choose $\sigma=\varepsilon/(2L)$ to obtain $\mathrm{KL}(\pi^{*}\|\pi_{\Theta})\leq\varepsilon$. Finally, by Lemma 2 (or the finiteness of $\mathcal{D}$), the achievable $\sigma$ tends to 0 as width $W\to\infty$, proving $\lim_{W\to\infty}\mathcal{E}(\mathcal{F}_{\mathrm{MLP}})=0$. ∎
Proposition 1.
Let $\mathcal{G}_{m}(\alpha,\beta)$, $\mathcal{F}_{m}(\alpha,\beta)$, and $\mathcal{C}_{m}^{k}(\alpha,\beta)$ be as in Definitions 12. For $n\in\mathbb{N}_{>0}$, let the large-generator class be
$\mathcal{G}_{M}(\alpha,\beta,n)\ :=\ \bigl\{\,g_{M}\ \bigm|\ W(g_{M})\ \geq\ k\alpha+n,\ \ D(g_{M})\ \geq\ \beta\,\bigr\},$
with induced policy class $\mathcal{F}_{M}(\alpha,\beta,n):=\{\mathrm{softmax}\!\circ g_{M}:\ g_{M}\in\mathcal{G}_{M}(\alpha,\beta,n)\}$. Assume the list-generation domain has a finite effective index set $\mathcal{D}=\{(i,l_{<i},j):j\in\mathcal{A}(i,l_{<i})\}$ and the activation $\sigma$ enjoys a universal-approximation property on compact sets (e.g., standard MLP activations). Then
$\mathcal{C}_{m}^{k}(\alpha,\beta)\ \subseteq\ \overline{\mathcal{F}_{M}(\alpha,\beta,n)},$
where the closure is taken w.r.t. uniform convergence of masked conditionals on $\mathcal{D}$.
Proof.
Fix any mixture element $\pi_{\mathrm{mix}}\in\mathcal{C}_{m}^{k}(\alpha,\beta)$. By Definition 2, there exist generators $g_{m,r}\in\mathcal{G}_{m}(\alpha,\beta)$, $r=1,\dots,k$, with logits $z_{r}(i,l_{<i},\cdot)$ on $\mathcal{A}(i,l_{<i})$ and mixture weights $\omega_{r}(i,l_{<i})\geq 0$ with $\sum_{r=1}^{k}\omega_{r}(i,l_{<i})=1$ such that
$\pi_{\mathrm{mix}}(\cdot\mid l_{<i})=\sum_{r=1}^{k}\omega_{r}(i,l_{<i})\,\mathrm{softmax}_{\mathcal{A}(i,l_{<i})}\!\bigl(z_{r}(i,l_{<i},\cdot)\bigr).$
Note that here we extend the mixture weights to be prefix-dependent, i.e., $\omega_{r}(i,l_{<i})$. This extension makes the proposition both stricter and more flexible, thereby broadening its generalization capability. If one wishes to exactly follow Definition 2, the weights can simply be degenerated to $\omega_{r}(l)$.
Step 1 (block-diagonal embedding of the $k$ small generators). Without loss of generality, pad each $g_{m,r}$ to width exactly $\alpha$ per hidden layer by adding zero weights/units. Construct a depth-$\beta$, width-$M$ network $g_{M}$ with $M\geq k\alpha+n$ whose hidden layers are partitioned into $k$ disjoint generator blocks of width $\alpha$ and one evaluator block of width $n$:
$M\;=\;\underbrace{\alpha+\cdots+\alpha}_{k\ \text{blocks}}\ +\ \underbrace{n}_{\text{evaluator}}.$
For layers $1,\dots,\beta-1$, set the large-layer weights to be block-diagonal so that the $r$-th generator block exactly replicates the corresponding layer of $g_{m,r}$, and the evaluator block either copies its previous state or computes auxiliary features (details in Step 2). Thus, after $\beta-1$ hidden layers, the first $k\alpha$ coordinates of the big network’s hidden state equal the concatenation of the $\beta-1$-th hidden activations of $\{g_{m,r}\}_{r=1}^{k}$.
Step 2 (parameterizing the evaluator weights $\omega$ with k1k-1 degrees of freedom). Use the evaluator block (of width $n\geq k-1$) to produce mixture logits $u(i,l_{<i})\in\mathbb{R}^{k}$ with the constraint that one coordinate is fixed as a reference (e.g., $u_{k}\equiv 0$), and define
$\omega_{r}(i,l_{<i})\ :=\ \frac{e^{u_{r}(i,l_{<i})}}{\sum_{q=1}^{k}e^{u_{q}(i,l_{<i})}}\quad(r=1,\dots,k),$
which realizes an arbitrary point in $\Delta^{k-1}$ through a k1k-1-dimensional parameterization. Because $\mathcal{D}$ is finite, the map $(i,l_{<i})\mapsto\omega(i,l_{<i})$ can be approximated arbitrarily well by the evaluator block via UAT.
Step 3 (combiner at depth $\beta$: realizing the log-sum-exp logits). Define for $j\in\mathcal{A}(i,l_{<i})$ the target combined logits
$\tilde{z}_{j}(i,l_{<i}):=\log\!\sum_{r=1}^{k}\frac{\omega_{r}(i,l_{<i})}{Z_{r}(i,l_{<i})}\,e^{z_{r,j}(i,l_{<i})},\qquad Z_{r}(i,l_{<i}):=\sum_{t\in\mathcal{A}(i,l_{<i})}e^{z_{r,t}(i,l_{<i})}.$
At the last hidden layer (the $\beta$-th nonlinear layer), allow cross-block connections from all generator blocks and the evaluator block into a width-$M$ hidden layer that serves as a single-hidden-layer approximator for the multivariate continuous mapping
$\Phi:\ \bigl(z_{1},\dots,z_{k},\omega\bigr)\ \mapsto\ \tilde{z}\quad\text{on the finite domain }\mathcal{D}.$
By universal approximation, there exist weights in this last hidden layer (and the final linear readout) so that the resulting $g_{M}$ satisfies
$\|\,g_{M}-\tilde{z}\,\|_{\infty;\mathcal{D}}\ \leq\ \sigma$
for any prescribed $\sigma>0$. Note that the depth requirement $D(g_{M})\geq\beta$ is met (we used exactly $\beta$ nonlinear layers), and the width requirement $W(g_{M})\geq k\alpha+n$ is used to house the $k$ embedded blocks ($k\alpha$ units) and the evaluator block ($n$ units).
Step 4 (from logits to conditionals). By the identity
$\mathrm{softmax}_{\mathcal{A}(i,l_{<i})}\!\bigl(\tilde{z}(i,l_{<i},\cdot)\bigr)\;=\;\sum_{r=1}^{k}\omega_{r}(i,l_{<i})\,\mathrm{softmax}_{\mathcal{A}(i,l_{<i})}\!\bigl(z_{r}(i,l_{<i},\cdot)\bigr),$
[…]
Remark 2 (Why $k\alpha+n$ and k1k-1 neurons for $\omega$).
The $k\alpha$ term guarantees disjoint capacity to exactly embed the $k$ small generators via block-diagonal copying across the first $\beta-1$ hidden layers. The additional $n$ units form an evaluator head; choosing $n\geq k-1$ suffices to parameterize the simplex $\Delta^{k-1}$ via softmax logits $u\in\mathbb{R}^{k}$ with one fixed reference coordinate, while also providing enough width for the last-layer universal approximation of the log-sum-exp combiner.
Proof of Theorem 1.
Step 1 (Coverage of $k$-mixtures by a single large generator). By Proposition 1,
$\mathcal{C}_{m}^{k}(\alpha,\beta)\ \subseteq\ \overline{\mathcal{F}_{M}(\alpha,\beta,n)},$
where the closure is taken w.r.t. uniform convergence of masked conditionals on the finite effective domain. Because $\pi^{*}(l)>0$ for all $l\in\mathcal{L}$ and $\mathcal{L}$ is finite, the map $\pi\mapsto\mathrm{KL}(\pi^{*}\|\pi)$ is continuous under uniform convergence of the conditionals. Hence, for every $n$,
$\mathcal{E}\bigl(\mathcal{F}_{M}(\alpha,\beta,n)\bigr)\ \leq\ \mathcal{E}\bigl(\mathcal{C}_{m}^{k}(\alpha,\beta)\bigr).$ (9)
Step 2 (Arbitrary accuracy by increasing width). By Theorem 3 (UAT-backed policy approximation), for every $\varepsilon>0$ there exists a fixed depth $L_{0}$ and a width threshold $W(\varepsilon,N,L)$, together with parameters $\Theta$, such that the induced policy $\pi_{\Theta}$ satisfies $\mathrm{KL}(\pi^{*}\|\pi_{\Theta})\leq\varepsilon$. Choose the fixed depth in Theorem 3 so that $L_{0}\geq\beta$, which is allowed by the theorem. Then, taking $n$ large enough to ensure $k\alpha+n\geq W(\varepsilon,N,L)$, we have $\pi_{\Theta}\in\mathcal{F}_{M}(\alpha,\beta,n)$ and therefore
$\mathcal{E}\bigl(\mathcal{F}_{M}(\alpha,\beta,n)\bigr)\ \leq\ \varepsilon.$ (10)
Since $\varepsilon>0$ was arbitrary, it follows that $\lim_{n\to\infty}\mathcal{E}(\mathcal{F}_{M}(\alpha,\beta,n))=0$.
Step 3 (Strict improvement over the $k$-mixture space). By Theorem 2, there exists $\delta>0$ such that $\mathcal{E}(\mathcal{C}_{m}^{k}(\alpha,\beta))=\delta$. Apply Step 2 with $\varepsilon:=\delta/2$. Then for some $n_{0}$,
$\mathcal{E}\bigl(\mathcal{F}_{M}(\alpha,\beta,n_{0})\bigr)\ \leq\ \delta/2\ <\ \delta\ =\ \mathcal{E}\bigl(\mathcal{C}_{m}^{k}(\alpha,\beta)\bigr).$
By monotonicity in $n$ (the class $\mathcal{F}_{M}(\alpha,\beta,n)$ enlarges with $n$), the strict inequality holds for all $n\geq n_{0}$. Combining with equation 9 concludes the proof of both statements. ∎
Appendix B Reward Model Training